OpenStudy (anonymous):

pleeeease help me!! tan x=7/24 , x Є [π,3π/2] cos x =?? sin x=??

OpenStudy (mathteacher1729):

Do you know how to draw the reference triangle for tan(x) = 7/24 ?

OpenStudy (anonymous):

no

OpenStudy (mathteacher1729):



OpenStudy (mathteacher1729):

Do you know how to find the length of the hypotenuse?

OpenStudy (anonymous):

no x|x

OpenStudy (mathteacher1729):

Have you heard of the phythagorean theorem?

OpenStudy (anonymous):

ooh yeah okay i forgot that way sqrt{24^2+7^2}

OpenStudy (mathteacher1729):

Yup yup. :) So what is \[\sqrt{24^2+7^2}\]

OpenStudy (anonymous):

sqrt{24^2+7^2} = 25 so cos x = 24/25 sin x = 7/25

OpenStudy (mathteacher1729):

You got it! :D

OpenStudy (anonymous):

thaanks a loot !!

OpenStudy (anonymous):

but cos x = 24/25 is not in the interval .. what shall i do for that

OpenStudy (anonymous):

@mathteacher1729